OpenStudy (anonymous):

what's the domain and the range of f(x)= square root of 4-x^2

OpenStudy (anonymous):

under radical must be non negative

OpenStudy (anonymous):

\[f(x)=\sqrt{4-x^2}\]

OpenStudy (anonymous):

4 - x^2 > = 0

OpenStudy (anonymous):

Solve for x

OpenStudy (anonymous):

so the answer would be x=2 or -2 ??????

OpenStudy (anonymous):

So.... Domain = [ -2 ,2]

OpenStudy (anonymous):

@mukushla am i correct..))

OpenStudy (anonymous):

yes u are... fatemah do u know why? D=[-2,2]

OpenStudy (anonymous):

yes because less or more than these values will give us negative values which cant be put under an even root :) so whats the value of the range??

OpenStudy (anonymous):

ok :) make 'x' ur subject..and of course we know domain of inverse function is range of function itself

OpenStudy (anonymous):

\[y=\sqrt{4-x^2}\]there is a restriction here y>=0 solve for x\[y^2=4-x^2\]\[x^2=4-y^2\]\[x=\sqrt{4-y^2}\]

OpenStudy (anonymous):

so the range is also = [-2,2] ???

OpenStudy (anonymous):

we had a restriction before doing inverse stuff\[y=\sqrt{4-x^2}\]this implies that \[y \ge0\]

OpenStudy (anonymous):

so whats the value of the range??

OpenStudy (anonymous):

[0,2] why?

OpenStudy (anonymous):

so the values of the domain will give us 0 , 1 or 2 ????

OpenStudy (anonymous):

every number between 0 and 2 :)